OpenStudy (anonymous):

Suppose a parabola has vertex (–4, 7) and also passes through the point (–3, 8). How do I write an equation for this is vertex form?

OpenStudy (amistre64):

start with the vertex form with a scalar y = m(x-h)^2 + k input the other point to determine the value of m b = m(a-h)^2 + k (b-k)/(a-h)^2 = m